How many grams are there in 7.40 moles of AGNO3
1
Expert's answer
2020-05-19T08:43:42-0400

Solution.

"n = \frac{m}{M}"

"m = n \times M"

"m = 7.40 \times 169.87 = 1257.04 \ g"

Answer:

"m = 1257.04 \ g"